# Responsible forecasting: identifying and typifying forecasting harms

### v1

@@ Abstract
Data-driven organizations around the world routinely use forecasting methods to improve their planning and decision-making capabilities. Although much research exists on the harms resulting from traditional machine learning applications, little has specifically focused on the ethical impact of time series forecasting. Yet forecasting raises unique ethical issues due to the way it is used in different organizational contexts, supports different goals, and involves different data processing, model development, and evaluation pipelines. These differences make it difficult to apply machine learning harm taxonomies to common forecasting contexts. We leverage multiple interviews with expert industry practitioners and academic researchers to remedy this knowledge gap by cataloguing and analysing under-explored domains, applications, and scenarios where forecasting may cause harm, with the goal of developing a novel taxonomy of forecasting-specific harms. Inspired by Microsoft Azure taxonomy for responsible innovation, we combined a human-led inductive coding scheme with an AI-driven analysis centered on the extraction of key taxonomies of harm in forecasting. The taxonomy is designed to guide researchers and practitioners and encourage ethical reflection on the impact of their decisions during the forecasting process. A secondary objective is to create a research agenda focused on possible forecasting-related measures to mitigate harm. Our work extends the growing literature on machine learning harms by identifying unique forms of harm that may occur in forecasting.
00footnotetext: Corresponding author: Bahman Rostami-Tabar, email: rostami-tabarb@cardiff.ac.uk
Keywords: time series forecasting, harm, models, algorithms, AI harms, ethics, data science
@@ 1 Introduction
But generating high-quality forecasts is difficult. Making matters worse, the probabilistic nature of forecasting can be difficult to square with our traditional legal and moral intuitions around blame and responsibility for harm. Consider the earthquake that occurred in the town of L’Aquila, Italy in 2009111https://www.theguardian.com/science/2012/oct/22/scientists-convicted-manslaughter-earthquake, which killed approximately 300 residents. A local jury convicted a group of seismologists of manslaughter for failing to warn the town of the earthquake. Should the seismologists be held morally and legally responsible for the deaths? The L’Aquila case highlights the ethical complexities facing various actors and institutions whose work involves producing or publishing forecasts.
@@
Despite the pervasive use of forecasting in business and society, we know surprisingly little about the harms caused during the practice of forecasting. There are several potential reasons for this. First, forecasting often takes place behind closed doors, thus making it difficult for external researchers to study. Second, forecasting—in contrast to traditional supervised learning with cross-sectional data—generally focuses on predicting the behavior of aggregates rather than individual units, making it hard to identify harms to specific individuals or organizations. The focus on aggregates also means that personal data generally plays a smaller role in forecasting. Third, unlike some applications of AI and machine learning that have provoked major media and journalism reporting (e.g., facial recognition or criminal risk predictions), applications of forecasting have tended to receive much less popular coverage. This last reason is perhaps again a symptom of the fact that much forecasting is done within organizations.
The relative neglect of research on forecasting harms means that we do not have a clear understanding of the factors that drive forecasting failures. We do not know how and which harms result from specific failures in the forecasting workflow. Even worse, many organizations using forecasting techniques lack incentives for publicly reporting forecasting failures and resulting harms, and some may even intentionally hide their failures to maintain a reputation of expertise (Bogdanich & Forsythe 2023). Beyond the algorithmic opacity intrinsic to the forecasting model (Burrell 2016), organizational opacity adds a layer of complexity to the study of forecasting harms. For example, the “problem of many hands” describes how bureaucratic complexity makes it difficult to attribute responsibility for harms to specific units or individuals (Nissenbaum 1996).
Forecasting failures have many possible root causes depending on where they occur in the forecasting process. Organizations may rely on forecasts despite lacking high-quality data or the requisite data science expertise. In other cases, failures may result from a mismatch of forecasting metrics and methods with organizational goals. Yet other failures stem from well-intentioned actors who misguidedly adjust undesirable forecasts. Adjusting forecasts may not only reduce forecast accuracy, but it can lead to other—often difficult to quantify—costs in terms human lives, the environment, people’s convenience and, ultimately, our trust in forecasting. For instance, the result of the 2016 US presidential election raised doubts in the minds of many citizens and political analysts about the validity of forecasting methods (Gelman 2021). As one prominent financial modeler bluntly put it, “some forecasters cause more damage to society than criminals” (Nassim 2007).
@@ 2.1 The concept of harm
Recognizing a forecasting harm requires a theory of harm to guide us and decide questions of moral and legal responsibility for harms. Yet, as with many concepts that have ethical and political implications, the concept of harm is broad and contested, often invoking related notions of wrongs, justice, duties, rights, responsibility, interests, intentions, causality, care, liability, and consent, to name just a few (Feinberg 1984). Our goal is not to articulate a novel theory of harm specific to forecasting, but to leverage the pre-existing harm-related concepts and frameworks developed and refined by legal scholars, philosophers, and (increasingly) AI ethics researchers.
Nearly all human cultures have subscribed to some version of the idea that we should, where reasonable, avoid doing harm to others (Bradley 2012). The prohibition on harm is so important that it is presumed to also apply to intelligent machines created by humans. For instance, the first of Isaac Asimov’s influential “three laws of intelligent machines” begins with the command: “A robot may not injure a human being, or, through inaction, allow a human being to come to harm” (Asimov 1978). So while there is nearly universal consensus that intelligent machines should avoid inflicting harm on humans, a more difficult question is what constitutes a harm. Here, we turn to fields that specialize in such definitional questions: philosophy and law.
Philosophical and legal accounts generally focus on three senses of harm: 1) something being damaged or broken; 2) unjust or wrongful treatment; and 3) the “defeating of an interest” (Feinberg 1984, pg. 33). For our purposes, we treat the concept of harm as a cluster concept, and provisionally define harm as unjustly or wrongly setting back an interest of an individual or collective of individuals (e.g., an organization).222 It is worth noting—especially since the AI ethics literature often invokes human rights (Kriebitz & Lütge 2020, Aizenberg & Van Den Hoven 2020)—that rights function as legal devices to protect an entity’s interests and imply (sometimes burdensome) correlative duties on others (O’Neill 2002). Legal documents that draw on fundamental rights, such as the EU’s GDPR, commonly rely on “balancing tests” to consider how the exercise of a data subject’s right justifies imposing a correlative duty on another entity (i.e., an organization wishing to process such data for business reasons) (Greene et al. 2019).
@@ 2.2 Typologies of AI harms
Technological harms can arise at multiple levels of society. Shelby et al. (2023) constructed a taxonomy of sociotechnical harms at micro-, meso- and macro-levels stemming from the use of algorithms. They thematically list the harms as representational harms, allocative harms, quality of service harms, interpersonal harms, and social system harms. Broadly, these harms relate to how social groups are represented in inputs and outputs, how these representations influence resource distribution decisions, how optimization choices relate to model performance for various users, how the model outputs affect social relations and properties of social systems, such as their level of stability or equality. Perhaps due to forecasting’s traditional focus on aggregate behavior, allocative and social systems harms are the most relevant to this discussion, while representational harms relating to the reinforcement of undesirable social and cultural stereotypes and beliefs are less relevant (see e.g., Barocas et al. (2017), Suresh & Guttag (2021)).
@@
As with many new technologies that are rapidly introduced into society, newly identified harms can ignite streams of research. For instance, the rise of deep neural networks and large labeled image datasets such as ImageNet, stimulated a wave of research into how various morally-charged labels were associated with certain demographic groups (Denton et al. 2021). Accordingly, with the rise of large language models (LLMs), a growing body of research is focuses on harms specific to LLMs. A number of LLM harms have been documented relating to discrimination, exclusion and toxicity, human-computer interaction harms, and automation, access, and environmental harms (Weidinger et al. 2022).
@@
Industry researchers at technology companies are the often the first to study the harms from new AI and ML-based technologies. A white paper from Microsoft (2023), which we refer to as the Microsoft Azure framework, specifically addresses responsible innovation and technology harm mitigation strategies. The framework classifies technology harms into four types: risk of injury, denial of consequential services, infringement on human rights, and the erosion of social and democratic structures. Risk of injury refers to physical injury due to, for example, over-reliance on safety features, inadequate fail-safes, and exposure to unhealthy agents during technology manufacturing or disposal. It also covers emotional or psychological injury due to emotional distress, distortion of reality, reduced self-esteem, addiction, identity theft, and misattribution. Denial of consequential services reflects opportunity loss such as employment, housing, insurance, and educational discrimination, as well as digital divide and loss of choice. Technologies that automate decisions can lead to credit discrimination, differential pricing, economic exploitation, and devaluation of individual expertise. Infringement of human rights involves privacy, liberty, and dignity harms via dehumanization (Aizenberg & Van Den Hoven 2020) and changes in how people perceive and engage with each other in society. The erosion of social and democratic structures includes harms related to amplification of power inequality, behavioral manipulation and deception, and stereotype reinforcement.
@@ 2.3 From harms in AI to harms in forecasting
Research on responsible AI and machine learning is now a mainstream research topic (Dignum 2019, Barocas et al. 2023). Yet little research has focused on how these ideas can be applied to forecasting pipelines, especially where forecasts influence human behavior and social systems. [^3] We distinguish forecasting from other similar terms using the following criteria: 1) forecasting is forward-looking, providing estimations of the variable of interest’s value at one or more future time periods; 2) forecasting is based on historical observations of the variable of interest (i.e., time series); and 3) the forecasting models can include past and future values of other exogenous variables (Hewamalage et al. 2021, Makridakis et al. 2020). Hobday et al. (2019) examines ethical issues stemming from forecasting in the context of marine ecology, and proposes ten guiding principles for ethical forecasting, including transparency, accurate uncertainty representation, stakeholder education, and performance reviews. These principles, although drawn from the specific domain of marine ecology, broadly align with the harm mitigation strategies we describe in Section 5.3.
@@ 2. Produce forecasting model cards
Forecasters should ensure that forecast users and consumers—the stakeholders— know what the forecast represents to prevent potential manipulation or misinterpretation. Forecast stakeholders should have access to information about the data, models, assumptions, the conditions under which the forecast applies, and forecast variability. Such transparency enables stakeholders to properly interpret and trust the forecasts. Model cards (Mitchell et al. 2019) could a viable solution to promote more standardized forecast reporting. Standardized reporting is used in clinical trials to summarize drug efficacy data and convey the uncertainty attached to regulatory decisions (Fischhoff & Davis 2014). Forecast model cards could provide a standardized disclosure mechanism, not unlike food nutrition labels or publicly available financial reports for firm investors, that help forecast stakeholders grasp the limitations and potential biases of the forecasts produced by a forecasting model. For instance, model cards could help guard against the file-drawer problem in forecasting, and mitigate the “problem of many hands” by encouraging accountability for forecasts by giving the name of an accountable individual or organization to be contacted in case of concern. Further, publicly available forecasting model information would allow stakeholders to engage in more informed decisions about where forecasting should and should not be applied, given the level of forecasting uncertainty and the costs of error.
@@ 7 Conclusions
One related question concerns responsibility for the harms caused by forecasting. Whether a poorly produced forecast actually results in harm can depend on luck. This complexity may spur discussion within the forecasting community around the adequacy of existing ethical guidelines or help inspire new “standards of care” uniquely adapted to forecasting. A greater appreciation and recognition of the harms of forecasting may even motivate steps towards the professional licensing of forecasters, particularly those working in high-risk domains or during emergency periods such as a global pandemic. For example, during the COVID-19 pandemic in Australia, the delay in publishing forecasts led to public confusion and poor decision-making, underscoring the potential harm of withholding critical predictive information. Indeed, professionalization could represent a broader, systemic approach to harm reduction that, for instance, involves institutionalizing some of the harm mitigation strategies discussed in Section 5.3.
@@
Acknowledgments
We would like to express our gratitude to the Notre Dame-IBM Technology Ethics Lab for their generous funding and support of this project, Project ID 522310. We also thank Dr. Nathaniel Raymond for his initial engagement during the early phases of the funding application. Fifi Ding assisted with the AI-driven analysis. Rob Hyndman is a member of the Australian Research Council Industrial Transformation Training Centre in Optimisation Technologies, Integrated Methodologies, and Applications (OPTIMA), Project ID IC200100009. Galit Shmueli’s research is partially supported by Taiwan National Science & Technology Council, research grant 111-2410-H-007-030-MY3.

### v3

# Good intentions, unintended consequences: exploring forecasting harms

@@ Abstract
Organizations worldwide that rely on data-driven approaches regularly employ forecasting methods to enhance their planning and decision-making processes. While extensive research has examined the harms associated with traditional machine learning applications, relatively little attention has been given to the ethical implications of time series forecasting. However, forecasting presents distinct ethical challenges due to its diverse organizational applications, varied objectives, and unique data processing, model development, and evaluation workflows. These distinctions complicate the direct application of existing machine learning harm taxonomies to common forecasting scenarios. To address this gap, we conduct multiple interviews with industry experts and academic researchers, systematically identifying and analyzing under-explored domains, use cases, and potential risks associated with forecasting. Our objective is to develop a novel taxonomy of forecasting-specific harms. Drawing inspiration from Microsoft Azure’s taxonomy for responsible innovation, we integrate a human-led inductive coding approach with AI-driven analysis to extract key categories of harm in forecasting. This taxonomy aims to support researchers and practitioners by fostering ethical reflection on their decision-making throughout the forecasting process. Additionally, we seek to establish a research agenda focused on identifying measures to mitigate potential harms in forecasting. By highlighting unique risks within forecasting, our work contributes to the broader discourse on machine learning ethics.
Keywords: forecasting, harm, models, algorithms, ethics, data science
@@ 1 Introduction
Despite the pervasive use and benefits of forecasting in business and society, we know surprisingly little about the harms specific to forecasting. Even popular textbooks on forecasting (Montgomery et al. 2015, Hyndman & Athanasopoulos 2021) do not substantively engage with the issue of harm [^1] The single mention of “harm” we found comes from the 600-plus-page textbook Montgomery et al. (2015, pg. 21), who asks the reader to discuss “the harm or penalty associated with a bad forecast” as a homework exercise. or discuss responsible forecasting practices. Very little work exists on these topics, yet in our experience most forecasters see value in being more transparent about its potential harms on individuals and society. This apparent tension piqued our curiosity.
Indeed, our interest in the topic of forecasting harm gradually emerged from candid discussions with forecasting experts in both industry and academia. Forecasters frequently lamented the lack of awareness and the need for a structured analysis of forecasting’s potential harm. For instance, the Editor-in-Chief of Foresight at the International Institute of Forecasters described the topic of forecasting harm as an “unexplored area,” stating, “I hadn’t heard this question posed like this before, but just bringing awareness makes a difference.” He emphasized that identifying common ways forecasting introduces harm could itself help mitigate negative effects. Further insights emerged from a panel on ethical considerations in forecasting organised by the lead author, featuring experts from the International Organization for Migration, UNHCR Innovation, the Neukom Institute for Computational Science, and Yale University. Panellists discussed forecasting’s unintended harms in migration, refugee crises, and public health, emphasizing the need for research and guidelines to mitigate negative consequences. They underscored how forecasting errors can have real-world impacts, affecting vulnerable populations and leading to resource misallocation. Industry case studies further reinforced these concerns around the social impact of forecasting—a Google data scientist highlighted how flawed forecasts contributed to Target Canada’s collapse, resulting in a \$2.1 billion loss and the closure of 133 stores. Similar concerns were echoed in our interviews of 21 experts from academia and industry. One respondent noted, “I’ve seen people so fascinated by producing forecasts that they don’t necessarily think about where it will be used, either well or poorly, for good or for harm” (P21). Another participant emphasized the importance of creating awareness: “Rather than accepting a set of signals, we should pause and think about how we consume this information” (P18). Overall, our discussions with forecasting experts in industry and academic shaped our research agenda by highlighting a few key gaps: the lack of a systematic treatment of forecasting harms, the tangible consequences of using forecasts across domains, and the need for ethical guidelines to mitigate the risks of harm.
@@
What might explain these gaps in our understanding of forecasting harms? We offer several tentative hypotheses. First, forecasting often takes place behind closed doors, making it difficult for researchers external to an organization to study. Second, forecasting—in contrast to traditional supervised learning with cross-sectional data—generally focuses on predicting the behavior of aggregates rather than individual units, making it hard to identify harms to specific individuals or organizations. When harms to individuals are thinly spread out over society, no one single person is motivated enough to speak out. Also, the focus on aggregates also means that personal data generally plays a smaller role in forecasting, which may contribute to less legal and regulatory scrutiny. Third, unlike some applications of AI and machine learning that have provoked major media and journalism reporting (e.g., social credit scoring, facial recognition, and criminal risk predictions), applications of forecasting have tended to receive much less popular coverage. This last reason is perhaps again because much information about forecasting stays within organizations, making it difficult for an external individual or group to contest the quality or accuracy of a forecast.
@@
Note that for the purpose of this study we distinguish forecasting from other similar terms, such as “predictive AI” or “predictive analytics,” [^2] In practice many of the same statistical forecasting methods are used atemporally to, for instance, classify objects, estimate past numeric values, or retrieve information (Barocas et al. 2023). using the following heuristic criteria: 1) forecasting is forward-looking, providing estimations of the variable of interest’s value at one or more future time periods; 2) forecasting is based on historical observations of the variable of interest (i.e., time series); and 3) the forecasting models can include past and future values of other exogenous variables (Hewamalage et al. 2021, Makridakis et al. 2020).
@@ 2.1 The concept of harm
Recognizing a forecasting harm requires a theory of harm to guide questions of moral and legal responsibility for harms. Yet, as with many concepts that have ethical and political implications, the concept of harm is interrelated with contested notions of wrongs, justice, duties, rights, responsibility, interests, intentions, causality, care, liability, and consent, to name just a few (Feinberg 1984). This makes it hard to define harm without begging the question. Our goal is not to articulate a novel theory of harm specific to forecasting, but to leverage the pre-existing harm-related concepts and frameworks developed and refined by legal scholars, philosophers, and (increasingly) AI ethics researchers.
Nearly all human cultures have subscribed to some version of the idea that we should, where reasonable, avoid doing harm to others (Bradley 2012). The prohibition on harm is so important that it is presumed to also apply to intelligent machines created by humans. For instance, the first of Isaac Asimov’s influential “three laws of intelligent machines” begins with the command: “A robot may not injure a human being, or, through inaction, allow a human being to come to harm” (Asimov 1978). So while there is nearly universal consensus that intelligent machines should avoid inflicting harm on humans, a more difficult question is what constitutes a harm. Here we turn to fields that specialize in examining definitional questions: philosophy and law.
Philosophical and legal accounts generally focus on three senses of harm: 1) something being damaged or broken; 2) unjust or wrongful treatment; and 3) the “defeating of an interest” (Feinberg 1984, pg. 33). In this initial study, we treat the concept of harm as a cluster concept, and provisionally define harm as unjustly or wrongly setting back an interest of an individual or collective of individuals (e.g., an organization). [^3] It is worth noting—especially since the AI ethics literature often invokes human rights (Aizenberg & Van Den Hoven 2020)—that rights function as legal devices to protect an entity’s interests and imply (sometimes burdensome) correlative duties on others (O’Neill 2002). Legal documents that draw on fundamental rights, such as the EU’s GDPR, commonly rely on “balancing tests” to consider how the exercise of a data subject’s right justifies imposing a correlative duty on another entity (i.e., an organization wishing to process such data for business reasons) (Greene et al. 2019). Future research might explore how to best define the normative terms “unjust” or “wrong” in the context of forecasting.
@@ 2.2 Typologies of AI harms
Technological harms can arise at multiple levels of society. Shelby et al. (2023) constructed a taxonomy of sociotechnical harms at micro-, meso- and macro-levels stemming from the use of algorithms. They thematically list the harms as representational harms, allocative harms, quality of service harms, interpersonal harms, and social system harms. Broadly, these harms relate to how social groups are represented in inputs and outputs, how these representations influence resource distribution decisions, how optimization choices relate to model performance for various users, how the model outputs affect social relations and properties of social systems, such as their level of stability or equality. Due to forecasting’s traditional focus on aggregate behavior, allocative and social systems harms appear the most relevant to forecasting, while representational harms (i.e., the reinforcement of undesirable social and cultural stereotypes and beliefs) seem less relevant (see e.g., Barocas et al. (2017), Suresh & Guttag (2021)).
@@
As with many new technologies that are rapidly introduced into society, newly identified harms can ignite streams of research. For instance, the rise of deep neural networks and large labeled image datasets such as ImageNet, stimulated a wave of research into how various morally-charged labels were associated with certain demographic groups (Denton et al. 2021). The rise of large language models (LLMs) has also led to a growing body of research focused on harms specific to LLMs. There are documented discrimination, exclusion and toxicity, and human-computer interaction harms, as well as automation, access, and environmental harms (Weidinger et al. 2022).
@@
Industry researchers are the often the first to study the harms from new AI technologies. A white paper from Microsoft, which we refer to as the Microsoft Azure framework (Microsoft 2023), specifically addresses responsible innovation and technology harm mitigation strategies. The framework classifies technology harms into four types: risk of injury, denial of consequential services, infringement on human rights, and the erosion of social and democratic structures. Risk of injury refers to physical injury due to, for example, over-reliance on safety features, inadequate fail-safes, and exposure to unhealthy agents during technology manufacturing or disposal. It also covers emotional or psychological injury due to emotional distress, distortion of reality, reduced self-esteem, addiction, identity theft, and misattribution. Denial of consequential services reflects opportunity loss such as employment, housing, insurance, and educational discrimination, as well as digital divide and loss of choice. Technologies that automate decisions can lead to credit discrimination, differential pricing, economic exploitation, and devaluation of individual expertise. Infringement of human rights involves privacy, liberty, and dignity harms via dehumanization and changes in how people perceive and engage with each other in society. The erosion of social and democratic structures includes harms related to amplification of power inequality, behavioral manipulation and deception, and stereotype reinforcement.
@@ 2.3 From harms in AI to harms in forecasting
While research on responsible AI and machine learning is growing (Dignum 2019, Barocas et al. 2023), translating and applying these ideas and methods to forecasting pipelines is not straightforward. There is a pressing need for such work given the scale of forecasting’s influence on human behavior and social systems. One notable exception is Hobday et al. (2019), who examines ethical issues stemming from forecasting in the context of marine ecology, and proposes ten guiding principles for ethical forecasting. He recommends transparency, accurate uncertainty representation, stakeholder education, and performance reviews. These principles, although drawn from the specific domain of marine ecology, broadly align with the harm mitigation strategies we describe in Section 5.3.
@@ 3.1 Data collection and analysis techniques
Ethical approval for this study was obtained from the Cardiff Business School Research Ethics Committee (Approval ID: 1727) on 19 June 2023. The study adhered to Cardiff University’s ethical principles as outlined in the Policy on the Ethical Conduct of Research Involving Human Participants, Human Material, or Human Data. Informed consent was obtained from all participants, data were anonymized after data collection, and measures were implemented to ensure confidentiality and data security.
@@ 2. Produce forecasting model cards
Forecasters should ensure that forecast consumers—the stakeholders— know what the forecast represents to prevent potential manipulation or misinterpretation. Forecast stakeholders should have access to information about the data, models, assumptions, the conditions under which the forecast applies, and forecast variability. Such transparency enables stakeholders to properly interpret and trust the forecasts. Model cards (Mitchell et al. 2019) could a viable solution to promote more standardized forecast reporting. Standardized reporting is used in clinical trials to summarize drug efficacy data and convey the uncertainty attached to regulatory decisions (Fischhoff & Davis 2014). Forecast model cards could provide a standardized disclosure mechanism, not unlike food nutrition labels or publicly available financial reports for firm investors, that help forecast stakeholders grasp the limitations and potential biases of the forecasts produced by a forecasting model. For instance, model cards could help guard against the file-drawer problem in forecasting, and mitigate the “problem of many hands” by encouraging accountability for forecasts by giving the name of an accountable individual or organization to be contacted in case of concern. Further, publicly available forecasting model information would allow stakeholders to engage in more informed decisions about where forecasting should and should not be applied, given the level of forecasting uncertainty and the costs of error.
@@ 7 Conclusions
One related question concerns responsibility for the harms caused by forecasting. Whether a poorly produced forecast actually results in harm can depend on luck. This complexity may spur discussion within the forecasting community around the adequacy of existing ethical guidelines or help inspire new “standards of care” uniquely adapted to forecasting. A greater appreciation and recognition of the harms of forecasting may even motivate steps towards the professional licensing of forecasters, particularly those working in high-risk domains or during emergency periods such as a global pandemic. Indeed, professionalization could represent a broader, systemic approach to harm reduction that, for instance, involves institutionalizing some of the harm mitigation strategies discussed in Section 5.3.
@@
Declaration of Conflicting Interests
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.